The sides of a triangle are in the ration 5 : 12 : 13. What is the length of each side of the triangle if the perimeter of the triangle is 15 in.?
Found 2 solutions by sim9739, checkley77:
Answer by sim9739(6)

as the ratio is 5:12:13,then we can let a variable n such that the triangle has the length of sides 5n, 12n and 13n,
15 = 5n+12n+13n
15 = 30n
thus, n = 0.5
so the length of each sides are 5*0.5, 12*0.5 and 13*0.5
they are 2.5, 6 and 6.5

Answer by checkley77(12844)

15/(5+12+13)15/30=.5 Which is the reduction factor for all sides. Thus:
5*.5+12*.5+13*.5=15
Proof:
2.5+6+6.5=15
15=15